# Capacity of entanglement for scalar fields in squeezed states

### v1

@@ 1 Introduction
where $\lambda_{\rm max}$ is the largest eigenvalue of the reduced density matrix and $\mathcal{N}$ denotes the number of nonvanishing eigenvalues of $\rho_{A}$. Besides the already mentioned case of the entanglement and Renyi entropies, there are many attempts to construct new information theoretic measures for studying the entanglement structure in more general setting, e.g., mutual information [8], logarithmic negativity [9] and entanglement of purification [10]. However, in this paper, we focus on another measure that has recently entered this discussion which is called the capacity of entanglement [11]
@@
Thus the capacity of entanglement characterizes the width of the eigenvalue spectrum of the reduced density matrix such that for a maximally entangled state, i.e., $\rho_{A}\propto I$, it vanishes. Indeed, in this case the Renyi entropies are independent of $n$ and we have a flat entanglement spectra. Recently, there have been many attempts to investigate various properties of this quantity in different setups both in the field theory and holography which have led to a remarkably rich and varied range of new insights, e.g., [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In particular, an interesting observation in [14] was that systems where are all entanglement is carried by EPR pairs have zero capacity of entanglement. Thus, whenever we find that $C_{E}\sim S_{E}$, e.g., $(1+1)$-dimensional CFTs at equilibrium, EPR pairs are not a very good approximation of the quantum state and randomly entangled pairs of qubits give a better description. Moreover, as proposed in this reference, restricting to field theories with holographic duals without higher derivative terms, the ratio $\frac{C_{E}}{S_{E}}$ turns out to be exactly equal to one. An interesting question is that whether this result also hold for more general QFTs without a holographic dual.
@@
Clearly, the capacity of entanglement depends on the choice of the original state of the system. The main aim of this article is to further investigate the state dependence of the capacity of entanglement. We study this dependence by considering a general class of Gaussian states which are solutions to the time dependent Schrödinger equation, the so-called squeezed states. As we will see the squeezed states are less classical than the vacuum state (which is a special case of a coherent state) and thus study the entanglement structure in a squeezed state may help us to gain a better insight into the quantum features of the system in question. In the following we will focus on a scalar field theory to address this problem. Nevertheless, having identified this particular QFT, we must first regulate the theory by placing it on a lattice, which reduces the system to an infinite family of quantum harmonic oscillators. Let us add that the entanglement entropy for this setup in squeezed states has been previously studied in [24]. These authors proposed an extension of the method introduced in [25, 26] to more general cases including squeezed states. Interestingly, using a systematic method they show that for a free scalar field theory in $(1+1)$-dimensions the entanglement entropy obeys the volume law rather than the area law in large squeezing limit. This is in contrast with the behavior of entanglement entropy in typical ground states which exhibits area law scaling where in $(1+1)$-dimensions is replaced with a logarithmic scaling. Further, their result is in agreement with the Page’s argument where for a typical pure quantum state of a joint system, the smaller subsystem is almost maximally mixed, showing little sign that the total system is pure [27]. Interestingly, the Page curve followed by the entropy of Hawking radiation [28]. In this context, an especially interesting question concerns how the scaling of the capacity of entanglement alters under a non-trivial squeezing when the total system is in a random squeezed state.
@@
Additionally, in [29, 30], using the holographic proposals, it was shown that the capacity of entanglement is a useful probe of the Hawking radiation during the black hole evaporation process. Interestingly enough, the authors argued that this measure shows a discontinuity or a peak at the Page time while the entanglement entropy varies smoothly. In this sense, one can consider $C_{E}$ as a convenient probe which can characterizes the order of a phase of the quantum systems. Indeed a primary motivation for this paper came from these efforts to study the behavior of different entanglement measures including the capacity of entanglement in random pure states and to compare our results to those obtained in the context of holography. Let us mention the caveat that this comparison involves two very different setups, namely a free quantum field theory with a single degree of freedom in our scalar field model versus a strongly coupled theory with a large number of degrees of freedom in holography. Thus there is no a priori reason to expect that the results should agree in the two cases.
@@ 2 Set-up
Indeed, they have the same minimal uncertainty value as found for the vacuum state. Further, the mean values of position and momentum operators in coherent state follow a classical orbit, e.g., $\langle\ddot{x}\rangle+\omega^{2}\langle x\rangle=0$, and hence these states correspond to the most classical states of the harmonic oscillator. It is also easy to show that any eigenstate of the annihilation operator, i.e., $a|\alpha\rangle=\alpha|\alpha\rangle$, where $\alpha$ is a complex parameter, represents a coherent state. Clearly, these states include the vacuum state as a special case with $\alpha=0$.
@@
$(\Delta x)_{\rm min}=\frac{1}{\sqrt{2m\omega e^{\pm z}}},\hskip 42.67912pt(\Delta p)_{\rm min}=\sqrt{\frac{m\omega e^{\pm z}}{2}},$ (2.9)
which shows the possibility of arbitrary compression of the position uncertainty at the expense of appropriate fluctuation in the momentum variable and vice-versa.
@@
In order to circumvent the divergences so as to obtain finite results for the measures, we must first regulate the theory by placing it on a lattice, which reduces the system to an infinite family of harmonic oscillators. Indeed, the general Hamiltonian for a system of $N$ coupled (one-dimensional) harmonic oscillators can be written as
@@
where $k_{1}$ determines the strength of the coupling between the two oscillators such that weak and strong coupling regimes correspond to $k_{1}\ll k_{0}$ and $k_{1}\gg k_{0}$ limits respectively. As it turns out, this simple model retains enough of the structure of the original problem that we will be able to learn several important lessons, which we can then carry over to the problem of computing the entanglement measures in our free scalar field theory. Let us recall that related investigations attempting to better understand the entanglement structure in general Gaussian states have also appeared in [31, 32, 33].
@@ 3 Preliminaries: $C_{E}$ for ground state
As a first step towards understanding different entanglement measures including capacity of entanglement in squeezed states, we would like to study the same quantities in a simpler set-up where the state is vacuum. We begin with the case of two coupled harmonic oscillators and then, having built up some intuition, we return to the field theory problem by generalizing these results to a lattice of coupled oscillators.
@@ 3.1 Two coupled harmonic oscillators
This recasts the problem as that of two decoupled oscillators, and thus it is now straightforward to solve for the eigenstates and eigenfrequencies of the Hamiltonian. In particular, we can write the vacuum state wave function as the product of the ground state wave functions for the two individual oscillators as follows
@@
Interestingly, we see that in this limit $C_{E}/S_{E}\ll 1$ which is consistent with the idea that the reduced density matrix becomes more and more maximally mixed as one increases the coupling.
@@
Figure 1: $S_{E}$ and $C_{E}$ (left), $S_{n}$ (middle) and $C_{n}$ (right) as functions of the coupling between the two harmonic oscillators.
We present the $n$-dependence of $S_{n}$ and $C_{n}$ for several values of the coupling in figure 2. Based on these plots, we see that the qualitative dependence of these measures on the coupling is similar to $n=1$ case. Also both measures decrease as we increase the Renyi index. Finally, as this simple example illustrates, the entanglement entropy and its fluctuations, characterized by the capacity of entanglement have no reason to be equal, except for a specific coupling. This result is different from what happens for QFTs with a holographic gravity dual (without higher derivative terms) where the ratio $C_{E}/S_{E}$ turns out to be exactly equal to one [14].
@@ 3.2 A lattice of oscillators and continuum limit
In this section we generalize our studies to a lattice of $N$ coupled harmonic oscillators, where the corresponding Hamiltonian is given by eq. (2.11), in specific directions. Drawing on our experience with the two coupled harmonic oscillators, this becomes a straightforward computation. In particular, as we saw above, the ground state wave function is described by a factorized Gaussian in the corresponding normal mode space. To see this we recall that as shown in [26] the matrix $K$ can be diagonalize by a real orthogonal similarity transformation, i.e., $K_{D}=UKU^{T}$. Then writing the Hamiltonian in terms of the normal modes reduces the problem to $N$ decoupled harmonic oscillators, which enables us to easily write the vacuum state wave function as
@@ 4 $C_{E}$ for harmonic systems in a squeezed state
In this section, we wish to return to our original problem which is computing the capacity of entanglement of a lattice of coupled oscillators in squeezed states. Again, as a warm up problem, we first consider the case of a single pair of coupled harmonic oscillators. Then, having built up some intuition, we shall generalize the problem in the next sections.
@@ 4.2 A chain of oscillators
The relation between the two spectra is $\xi_{j}=\frac{\tilde{\xi}_{j}-1}{\tilde{\xi}_{j}+1}$ where the normalization condition for the density matrix allows us to neglect $\tilde{\xi}_{j}<1$. Further, the eigenstates can be written in terms of the Hermite polynomials. We skip over the details of the calculation and we refer the interested reader to [24] for further details. Having the corresponding eigenvalues, the entanglement measures can be found using the same formulae as in the previous section, e.g., eq. (3.20). Now we are equipped with all we need to study the behavior of the entanglement spectrum and thereby other related quantities in a squeezed state. To do so, we employ a numerical treatment in the next section.
@@ 5.2 Squeezing all modes
where $\ell\equiv\widetilde{N}\epsilon$ and $L\equiv N\epsilon\rightarrow\infty$. Thus in this limit $\frac{\overline{S_{E}}}{\overline{C_{E}}}=z\gg 1$ which is a feature of a maximally mixed state. Moreover, the measures are both sensitive to the UV cutoff, but the ratio is finite. We conclude that at least for the free massless scalar field theory, the ratio of the leading terms is scheme independent. Of course, the numerical results depicted in the left panel for the mean entanglement entropy are consistent with the Page’s argument that in an arbitrary quantum state the entropy is close to maximal. Hence we expect entanglement entropy to scale with the volume (instead of the area) of the entangling region.
The right panel shows that this special behavior also holds for capacity of entanglement. In particular, we see that $\overline{C_{E}}$ is always continuous and varies smoothly. This feature contrasts with the corresponding results for the capacity of entanglement corresponding to the black hole evaporation process as previously noted in [29, 30]. Of course, this comparison involves two very different QFTs, namely a free quantum field theory which can be written as an infinite family of harmonic oscillators versus a strongly coupled theory with a large number of degrees of freedom in holography. Thus there is no a priori reason to expect that the results should agree in the two cases. Interestingly enough, based on eq. (5.1) we see that in large squeezing limit the relationship $C_{E}=S_{E}$ is completely broken. Indeed, as proposed in [14] perhaps such a relation in QFTs is a hint of a dual gravitational interpretation. Therefore, at least in this regime the corresponding squeezed states can not have a solution of a classical gravity theory as a holographic dual. This is the main clue that forbids us to compare our results with some previous studies on capacity of entanglement in the context of holography. Let us emphasize that we found similar results for $\overline{S_{n}}$ and $\overline{C_{n}}$ which are consistent with the asymptotic expansions we reported in the previous section, although we do not explicitly show the corresponding figures here.
@@ 6 Conclusions and discussions
We have also extended these studies to the system of $N$ coupled harmonic oscillators to investigate to what extent the squeezing parameter modifies the behavior of the entanglement measures including the capacity of entanglement. In particular, we have shown that where all modes lie in a squeezed state the measures are time-independent and proportional to the volume of the smaller subsystem in large squeezing limit. Again, we have found that in this limit the ratio of capacity over entropy is negligible and thus the reduced density matrix behaves as a maximally mixed state.
@@
Once again, considering the more realistic case where all modes lie in a squeezed state, the large $z$ behavior of the measures are completely independent of time. In particular, in this regime the scaling of the capacity of entanglement is not sensitive to the details of the time evolution of the reduced density matrix. We have also found numerically that at leading order $C_{E}$ is proportional to the volume of the smaller subsystem which is consistent with asymptotic scaling given by eq. (4.18). Interestingly, for an infinite chain of oscillators in the continuum limit the corresponding behaviors of $S_{E}$ and $C_{E}$ are given by eq. (5.1) which is a feature of a maximally mixed state. Here we recall that this equation holds in the large squeezing limit and thus $C_{E}/S_{E}\sim z^{-1}\rightarrow 0$. As we have mentioned before this means that the corresponding reduced density matrix can be approximated as proportional to the identity operator to the extent that its Renyi entropies are independent of $n$ and thus we have a flat entanglement spectra. Remarkably, in the holographic context one can produce the $n$-independent Renyi entropies by considering specific semiclassical states, the so-called fixed-area states [34, 35]. [^5] We thank Mohammad Hasan Vahidinia for bringing this to our attention and raising the following question. It is an interesting question whether or not a more concrete connection can be found between the squeezing states in the filed theory and fixed-area states of quantum gravity.
@@
To close our discussion, we would like to recall that our result for the capacity of entanglement for a family of random states, where $\overline{C_{E}}$ is continuous, contrasts with the holographic computations correspond to the black hole evaporation process reported in [29, 30]. Indeed, we expect that a meaningful comparison between the field theory and holographic results is achieved by considering strongly coupled theories with a large number of degrees of freedom. This investigation is beyond the scope of the present paper. It would be interesting to explore the behavior of $C_{E}$ in general setups including rational and holographic CFTs [36, 37, 38, 39]. Some additional topics to explore include generalizing our results to higher dimensions [40] or to non-relativistic models [41, 42]. Another interesting direction is to study the scaling of capacity of entanglement in theories which exhibit a volume law scaling of the entanglement entropy [43, 44, 45, 46]. In addition, we have been focused on the exploration of the capacity only in a free scalar theory, but it should be feasible to extend our work to more realistic cases with non-trivial interactions where we believe the capacity will exhibit more interesting features [47]. We leave the details of some interesting problems for future study [48].

### v3

@@ 1 Introduction
where $\lambda_{\rm max}$ is the largest eigenvalue of the reduced density matrix and $\mathcal{N}$ denotes the number of nonvanishing eigenvalues of $\rho_{A}$. Besides the already mentioned case of the entanglement and Renyi entropies, there are many attempts to construct new information theoretic measures for studying the entanglement structure in more general setting, e.g., mutual information [8], logarithmic negativity [9, 10] and entanglement of purification [11]. However, in this paper, we focus on another measure that has recently entered this discussion which is called the capacity of entanglement [12]
@@
Thus the capacity of entanglement characterizes the width of the eigenvalue spectrum of the reduced density matrix such that for a maximally entangled state, i.e., $\rho_{A}\propto I$, it vanishes. Indeed, in this case the Renyi entropies are independent of $n$ and we have a flat entanglement spectra. Recently, there have been many attempts to investigate various properties of this quantity in different setups both in the field theory and holography which have led to a remarkably rich and varied range of new insights, e.g., [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. In particular, an interesting observation in [15] was that systems where are all entanglement is carried by EPR pairs have zero capacity of entanglement. Thus, whenever we find that $C_{E}\sim S_{E}$, e.g., $(1+1)$-dimensional CFTs at equilibrium, EPR pairs are not a very good approximation of the quantum state and randomly entangled pairs of qubits give a better description. Moreover, as proposed in this reference, restricting to field theories with holographic duals without higher derivative terms, the ratio $\frac{C_{E}}{S_{E}}$ turns out to be exactly equal to one. An interesting question is whether this result also hold for more general QFTs without a holographic dual.
@@
Clearly, the capacity of entanglement depends on the choice of the original state of the system. The main aim of this article is to further investigate the state dependence of the capacity of entanglement. We study this dependence by considering a general class of Gaussian states which are solutions to the time dependent Schrödinger equation, the so-called squeezed states. As we will see the squeezed states are less classical than the vacuum state (which is a special case of a coherent state) and thus study the entanglement structure in a squeezed state may help us to gain a better insight into the quantum features of the system in question. In the following we will focus on a scalar field theory in $(1+1)$-dimensions to address this problem. Indeed, in order to avoid the ultraviolet divergences in the continuum limit, we should regulate the theory by placing it on a one dimensional spatial lattice. Thus the model reduces to an infinite chain of quantum harmonic oscillators. Let us add that the entanglement entropy for this setup in squeezed states has been previously studied in [26]. These authors proposed an extension of the method introduced in [27, 28] to more general cases including squeezed states. Interestingly, using a systematic method they show that for a free scalar field theory in $(1+1)$-dimensions the entanglement entropy obeys the volume law rather than the area law in large squeezing limit. This is in contrast with the behavior of entanglement entropy in typical ground states which exhibits area law scaling where in $(1+1)$-dimensions is replaced with a logarithmic scaling. Further, their result is in agreement with the Page’s argument where for a typical pure quantum state of a joint system, the smaller subsystem is almost maximally mixed, showing little sign that the total system is pure [29]. Interestingly, the Page curve followed by the entropy of Hawking radiation [30]. In this context, an especially interesting question concerns how the scaling of the capacity of entanglement alters under a non-trivial squeezing when the total system is in a random squeezed state.
@@
Additionally, in [31, 32], using the holographic proposals, it was shown that the capacity of entanglement is a useful probe of the Hawking radiation during the black hole evaporation process. Interestingly enough, the authors argued that this measure shows a discontinuity or a peak at the Page time while the entanglement entropy varies smoothly. In this sense, one can consider $C_{E}$ as a convenient probe which can characterizes the order of a phase of the quantum systems. Indeed a primary motivation for this work came from these efforts to study the behavior of different entanglement measures including the capacity of entanglement in random pure states and to compare our results to those obtained in the context of holography. Let us mention that this comparison involves two very different models, that is a free scalar theory with a single degree of freedom versus a holographic quantum field theory which is strongly coupled and has a large number of degrees of freedom.
@@ 2 Set-up
Indeed, they have the same minimal uncertainty value as found for the vacuum state. Further, the mean values of position and momentum operators in coherent state follow a classical orbit, e.g., $\langle\ddot{x}\rangle+\omega^{2}\langle x\rangle=0$, and hence these states correspond to the most classical states of the harmonic oscillator. It is also easy to show that any eigenstate of the annihilation operator, i.e., $a|\alpha\rangle=\alpha|\alpha\rangle$, where $\alpha$ is a complex parameter, represents a coherent state. Clearly, these states include the vacuum state as a special case with $\alpha=0$. Moreover, defining a unitary displacement operator $D(\alpha)=e^{\alpha a^{\dagger}-\alpha^{*}a}$, the coherent states can be generated from the vacuum, i.e., $|\alpha\rangle=D(\alpha)|0\rangle$.
@@
$(\Delta x)_{\rm min}=\frac{1}{\sqrt{2m\omega e^{z}}},\hskip 42.67912pt(\Delta p)_{\rm min}=\sqrt{\frac{m\omega e^{z}}{2}},$ (2.9)
which shows the possibility of arbitrary compression of the position uncertainty at the expense of appropriate fluctuation in the momentum variable and vice-versa. Interestingly, defining a unitary squeezing operator $S(\xi)=e^{\frac{1}{2}(\xi a^{\dagger 2}-\xi^{*}a^{2})}$, the squeezed states can be generated from the vacuum, i.e., $|\xi\rangle=S(\xi)|0\rangle$ where $\xi$ denotes a complex parameter. [^3] Note that this definition gives a squeezed vacuum state. In addition one can obtain a squeezed coherent state by acting the squeezing operator on a coherent state, i.e., $|\xi,\alpha\rangle=S(\xi)|\alpha\rangle=S(\xi)D(\alpha)|0\rangle$. For reviews on the main properties of coherent and squeezed states see e.g. Refs. [33, 34] and references therein.
@@
In order to circumvent the divergences so as to obtain finite results for the measures, we should regulate the above model by placing it on a lattice, which reduces the system to an infinite chain of coupled harmonic oscillators. Indeed, the general Hamiltonian for a system of $N$ coupled (one-dimensional) harmonic oscillators can be written as
@@
where $k_{1}$ determines the strength of the coupling between the two oscillators such that weak and strong coupling regimes correspond to $k_{1}\ll k_{0}$ and $k_{1}\gg k_{0}$ limits respectively. We see that this simple set-up maintains some interesting features of our original problem which helps us to better investigate the continuum scalar theory. Let us recall that related investigations attempting to better understand the entanglement structure in general Gaussian states have also appeared in [35, 36, 37].
@@ 3 Preliminaries: $C_{E}$ for ground state
As a first step toward understanding different entanglement measures including capacity of entanglement in squeezed states, we would like to study the same quantities in a simpler set-up where the state is vacuum. We begin with the case of two coupled harmonic oscillators and then, having built up some intuition, we return to the more involved problem by generalizing these results to a lattice of coupled oscillators.
@@ 3.1 Two coupled harmonic oscillators
Now we have two decoupled harmonic oscillators, and thus the vacuum state can be written as the product of the vacuum state wave functions for the two individual oscillators as follows
@@
Interestingly, we see that in this limit $C_{E}/S_{E}\ll 1$ which is consistent with the idea that the reduced density matrix becomes more and more maximally mixed as one increases the coupling. [^4] Note that as we have already mentioned, for a maximally mixed state $C_{E}$ exactly vanishes. Although, here we follow the terminology of [15] where the strength of the ratio between the capacity and entropy indicates how much a given reduced density matrix is close to a maximal distribution. Similar terminology used in [29] where the author encountered “almost” maximally mixed states. We would like to thank the anonymous referee for a useful comment on this point.
@@
Figure 1: $S_{E}$ and $C_{E}$ (left), $S_{n}$ (middle) and $C_{n}$ (right) as functions of the coupling between the two harmonic oscillators. The dashed lines indicate the large $n$ limit which are consistent with eq. (3.12).
We present the $n$-dependence of $S_{n}$ and $C_{n}$ for several values of the coupling in figure 2. Based on these plots, we see that the qualitative dependence of these measures on the coupling is similar to $n=1$ case. Also both measures decrease as we increase the Renyi index. Using eqs. (3.8) and (3.1) it is easy to find the large $n$ scaling of these measures as follows
$S_{n\rightarrow\infty}\sim-\log(1-\xi),\hskip 56.9055ptC_{n\rightarrow\infty}\sim\xi^{n}n^{2}(\log\xi)^{2},$ (3.12)
which shows that for any $10<\xi<$ 1, $C_{n\rightarrow\infty}$ vanishes. Indeed, our numerical results approach the above expressions in this limit as depicted in figure 1. Finally, as this simple example illustrates, the entanglement entropy and its fluctuations, characterized by the capacity of entanglement have no reason to be equal, except for a specific coupling. This result is different from what happens for QFTs with a holographic gravity dual (without higher derivative terms) where the ratio $C_{E}/S_{E}$ turns out to be exactly equal to one [15].
@@ 3.2 A lattice of oscillators and continuum limit
In this section we generalize our studies to a lattice of $N$ coupled harmonic oscillators, where the corresponding Hamiltonian is given by eq. (2.11), in specific directions. Similar to the previous case, the ground state wave function is described by a factorized Gaussian in the corresponding normal mode space. To see this we recall that as shown in [28] the matrix $K$ can be diagonalize by a real orthogonal similarity transformation, i.e., $K_{D}=UKU^{T}$. Then writing the Hamiltonian in terms of the normal modes reduces the problem to a chain of decoupled quantum harmonic oscillators. Thus the total vacuum state can be written as
@@ 4 $C_{E}$ for harmonic systems in a squeezed state
In this section, we wish to return to our original problem which is computing the capacity of entanglement of a chain of coupled oscillators in squeezed states. Again, as a warm up problem, we first consider the case of a single pair of coupled harmonic oscillators and then generalize the problem to more involved case consisting of $N$ degrees of freedom in the next sections.
@@ 4.2 A chain of oscillators
The relation between the two spectra is $\xi_{j}=\frac{\tilde{\xi}_{j}-1}{\tilde{\xi}_{j}+1}$ where the normalization condition for the density matrix allows us to neglect $\tilde{\xi}_{j}<1$. Further, the eigenstates can be written in terms of the Hermite polynomials. We skip over the details of the calculation and we refer the interested reader to [26] for further details. Having the corresponding eigenvalues, the entanglement measures can be found using the same formulas as in the previous section, e.g., eq. (3.21). Now we are equipped with all we need to study the behavior of the entanglement spectrum and thereby other related quantities in a squeezed state. To do so, we employ a numerical treatment in the next section.
@@ 5.2 Squeezing all modes
where $\ell\equiv\widetilde{N}\epsilon$ and $L\equiv N\epsilon\rightarrow\infty$. Thus in this limit $\frac{\overline{S_{E}}}{\overline{C_{E}}}=z\gg 1$ which is a feature of an “almost” maximally mixed state. Moreover, the measures are both sensitive to the UV cutoff, but the ratio is finite. We conclude that at least for the free massless scalar field theory, the ratio of the leading terms is scheme independent. Of course, the numerical results depicted in the left panel for the mean entanglement entropy are consistent with the Page’s argument that in an arbitrary quantum state the entropy is close to maximal. Hence we expect entanglement entropy to scale with the volume (instead of the area) of the entangling region.
The right panel shows that this special behavior also holds for capacity of entanglement. In particular, we see that $\overline{C_{E}}$ is always continuous and varies smoothly. This feature contrasts with the holographic results for the capacity of entanglement corresponding to the black hole evaporation process as previously noted in [31, 32]. Of course, their results should be hold in a strongly coupled theory with a large number of degrees of freedom which is completely different from our free scalar model. Thus there is no a priori reason to expect that the results should agree in these cases. Interestingly enough, based on eq. (5.1) we see that in large squeezing limit the relationship $C_{E}=S_{E}$ is completely broken. Indeed, as proposed in [15] perhaps such a relation in QFTs is a hint of a dual gravitational interpretation. Therefore, at least in this regime the corresponding squeezed states cannot have a solution of a classical gravity theory as a holographic dual. This is the main clue that forbids us to compare our results with some previous studies on capacity of entanglement in the context of holography. Let us emphasize that we found similar results for $\overline{S_{n}}$ and $\overline{C_{n}}$ which are consistent with the asymptotic expansions we reported in the previous section, although we do not explicitly show the corresponding figures here.
@@ 6 Conclusions and discussions
We have also extended these studies to the system of $N$ coupled harmonic oscillators to investigate to what extent the squeezing parameter modifies the behavior of the entanglement measures including the capacity of entanglement. In particular, we have shown that where all modes lie in a squeezed state the measures are time-independent and proportional to the volume of the smaller subsystem in large squeezing limit. Again, we have found that in this limit the ratio of capacity over entropy is negligible and thus the reduced density matrix behaves as an “almost” maximally mixed state.
@@
Once again, considering the more realistic case where all modes lie in a squeezed state, the large $z$ behavior of the measures are completely independent of time. In particular, in this regime the scaling of the capacity of entanglement is not sensitive to the details of the time evolution of the reduced density matrix. We have also found numerically that at leading order $C_{E}$ is proportional to the volume of the smaller subsystem which is consistent with asymptotic scaling given by eq. (4.18). Interestingly, for an infinite chain of oscillators in the continuum limit the corresponding behaviors of $S_{E}$ and $C_{E}$ are given by eq. (5.1) which shows that the reduced density matrix becomes more and more maximally mixed in the large squeezing limit such that $C_{E}/S_{E}\sim z^{-1}\rightarrow 0$. As we have mentioned before this means that the corresponding reduced density matrix can be approximated as proportional to the identity operator to the extent that its Renyi entropies are independent of $n$ and thus we have a flat entanglement spectra. Remarkably, in the holographic context one can produce the $n$-independent Renyi entropies by considering specific semiclassical states, the so-called fixed-area states [38, 39, 40, 41]. [^7] We thank Mohammad Hasan Vahidinia for bringing this to our attention and raising the following question. It is an interesting question whether or not a more concrete connection can be found between the squeezing states in the field theory and fixed-area states of quantum gravity.
@@
To close our discussion, we would like to recall that our result for the capacity of entanglement for a family of random states, where $\overline{C_{E}}$ is continuous, contrasts with the holographic computations correspond to the black hole evaporation process reported in [31, 32]. Indeed, we expect that a meaningful comparison between the field theory and holographic results is achieved by considering strongly coupled theories with a large number of degrees of freedom. This investigation is beyond the scope of the present paper. It would be interesting to explore the behavior of $C_{E}$ in general setups including rational and holographic CFTs [42, 43, 44, 45]. Some additional topics to explore include generalizing our results to higher dimensions [46] or to non-relativistic models [47, 48, 49, 50]. Another interesting direction is to study the scaling of capacity of entanglement in theories which exhibit a volume law scaling of the entanglement entropy [51, 52, 53, 54]. In addition, we have been focused on the exploration of the capacity only in a free scalar theory, but it should be feasible to extend our work to more realistic cases with non-trivial interactions where we believe the capacity will exhibit more interesting features [55]. We leave the details of some interesting problems for future study [56].